### v3

# LLM-Generated Heuristics for AI Planning: Do We Even Need Domain-Independence Anymore?

Abstract
Heuristics are a central component of classical AI planning, particularly in domain-independent settings where general applicability is prioritized over task-specific tuning. This work revisits that paradigm in light of recent advances in large language models (LLMs), which enable the automatic synthesis of heuristics directly from problem definitions—bypassing the need for handcrafted domain knowledge. We present a method that employs LLMs to generate instance-specific heuristic functions from planning tasks specified through successor generators, goal tests, and initial states written in a general-purpose programming language. These heuristics are compiled and integrated into standard heuristic search algorithms, such as greedy best-first search. Our approach achieves competitive, and in many cases state-of-the-art, performance across a broad range of established planning benchmarks. Moreover, it enables the solution of problems that are difficult to express in traditional formalisms, including those with complex numeric constraints or custom transition dynamics. We analyze the trade-offs between generality and performance, and argue that LLM-generated heuristics represent a promising direction for integrating machine learning with symbolic reasoning in planning.
Introduction
Deterministic AI planning, at its core, involves finding a sequence of actions that transitions an agent from an initial state to a goal state, given a formal description of the environment’s dynamics. To search effectively through large state spaces, planners rely on heuristics—functions that estimate the distance to the goal and guide the search toward promising regions.
Classical planning has traditionally focused on general, domain- and problem-independent heuristics, which offer broad applicability and avoid the need for task-specific design (Bonet and Geffner 2001). However, the performance of such heuristics can depend on the representation language and the structural characteristics of the domain. In more expressive settings, general heuristics may face limitations or require adaptation to remain effective in different versions of pddl (cf. Fox and Long 2003; Edelkamp and Hoffmann 2004; Gerevini and Long 2005).
In this work, we propose to automatically generate task-specific heuristic functions from formal problem definitions, using LLMs to generate task-specific heuristics, that are then used by a classic AI search algorithm. In this way, we benefit from the systematic and guaranteed search properties of the classic algorithm, while keeping the generation of the heuristics automatic, with no need for expert manual labor for heuristics design.
We represent each planning task using three components implemented in a general-purpose programming language: a successor generator, a goal test, and an initial state. These are provided as input to the LLM, which returns a heuristic function which can be tailored to the specific problem structure. This heuristic is compiled and used to guide search without further interaction with the LLM.
This approach has several advantages. First, it avoids repeated LLM calls during planning, greatly improving efficiency. Second, it enables the use of transparent, verifiable search components: both the LLM-generated heuristic and the underlying task specification are standard Rust code that can be debugged and tested. Third, it allows us to solve planning problems that involve constructs difficult to express in traditional frameworks—such as recursive goals, complex numeric conditions, or custom transition logic.
We show that LLMs can be used to improve general-purpose planners by using them to generate input-specific heuristics, rather than relying on general-purpose heuristics. We evaluate our approach on a set of standard AI planning benchmarks, translated into Rust representations suitable for LLM prompting. In many domains, the resulting heuristics outperform classical domain-independent methods and compete with state-of-the-art numeric planners. Additionally, our method is able to solve novel planning tasks that lack formal encoding in existing planning languages.
Related Work
This paper revisits AI planning by heuristic search, by replacing task-independent heuristics with LLM-generated task-specific heuristics. Below, we review related work on LLMs and planning. For a more comprehensive overview of LLMs as planning modelers, see the surveys by (Tantakoun, Zhu, and Muise 2025) and by (Aghzal et al. 2025).
Planning with LLMs
(Valmeekam et al. 2022) provides an important early evaluation of LLM-based planning, demonstrating that even on simple benchmarks, LLMs do not consistently exhibit reliable planning behavior. While it seems that LLMs cannot yet generate plans efficiently, our work shows that they can be used effectively as a component of a larger planning paradigm. Namely, LLMs can generate heuristics and search components rather than perform planning directly.
A closely related approach is presented by (Katz et al. 2024), who argue that existing LLM-based planning methods are neither sound nor complete and incur high computational costs due to repeated LLM calls. They propose an alternative, Thought of Search (ToS), in which LLMs are used to generate symbolic search components—successor functions and goal tests. These can then be executed efficiently without further model queries. Although their work focuses on small, finite search problems such as the 24 game and Mini Crosswords, they provide compelling evidence that such an approach can drastically reduce computational cost and improve correctness (solved by BFS). The work shows that with respect to planning their approach overperforms the planning directly with LLM approaches such as: Chain-of-Thought (Wei et al. 2022), ReAct (Yao et al. 2023b), ReWOO (Xu et al. 2023), Reasoning via Planning (Hao et al. 2023), Tree of Thoughts (Yao et al. 2023a), Graph of Thoughts (Besta et al. 2024), Reflexion (Shinn et al. 2023), Algorithm of Thought (Sel et al. 2024). The ToS approach runs a Breadth-First Search (BFS) using components—successor functions and goal tests generated by LLM. Our approach extends this intuition by employing heuristic search methods to address problems of significantly greater complexity than those solvable through naive BFS—specifically in AI planning, where transition systems often become infinite due to the presence of unbounded numeric fluents. Consequently, generating effective heuristics is particularly crucial, underscoring that the symbolic utilization of LLMs can be successfully applied beyond toy domains.
A recent study in a similar vein to our work is presented in (Corrêa, Pereira, and Seipp 2025). While their approach targets planning tasks specified in the Planning Domain Definition Language (PDDL), ours departs from this framework by operating directly on successor generators and goal-check functions. This shift affords greater representational flexibility, enabling the modeling of tasks that PDDL cannot express effectively. However, abandoning PDDL’s structured representation also limits our ability to exploit domain structure for heuristic derivation.
Heuristic Generation Outside of Planning
LLMs have also been explored for heuristic generation beyond automated planning, notably in domains such as online bin packing and TSP (Liu et al. 2024), as well as other combinatorial optimization problems (Ye et al. 2024). These approaches typically combine LLMs with evolutionary algorithms to evolve heuristics that guide the optimization process. However, due to fundamental differences in problem structure, these methods are not directly applicable to planning tasks without substantial adaptation.
Problem Generation
Recent work has explored using LLMs to generate domains and problem descriptions using pddl2.1 (Fox and Long 2003), among such works are (Guan et al. 2023; Liu et al. 2023; Oswald et al. 2024). For instance, (Silver et al. 2024) demonstrated that GPT-4 can output plans from parsed pddl2.1 inputs using simple, non-search-based strategies. (Guan et al. 2023) proposed an approach in which pddl2.1 actions are generated one at a time via LLM queries, refined using human feedback. A related concept appears in (Zhou et al. 2024), who use LLMs to generate, execute, and refine Python code for solving mathematical reasoning problems. Their findings support the idea that LLMs can produce verifiable code with minimal manual intervention. To reduce human involvement even further, we propose generating Rust code instead of Python. Since Rust is a compiled language with strong static typing, much of the burden of code verification can be offloaded to the compiler.
Code Generation Our approach relies on LLMs not to solve problems directly, but to generate search guidance in the form of state evaluation heuristic code. (Zhou et al. 2024) pursued a similar strategy in mathematical domains, showing that LLMs can generate and iteratively improve executable code. Our work builds on this insight, but targets search guidance functions within a planning framework. The viability of this approach is supported by recent advances in program synthesis and LLM-based code generation (Madaan et al. 2023; Zhang et al. 2023; Chen et al. 2024; Muennighoff et al. 2024; Zhong, Wang, and Shang 2024), which show that language models are increasingly capable of producing correct, type-safe, and semantically meaningful programs.
Deterministic AI Planning
AI planning refers to the problem of computing a sequence of actions that transforms a known initial state into one that satisfies a specified goal condition. Unlike reinforcement learning, which learns behavior through trial-and-error interaction with an environment, AI planning assumes full knowledge of the state transition dynamics and the goal. The planner searches the combinatorial space of states reachable via actions to construct a valid plan.
Formally, a deterministic planning problem can be defined as a tuple $\Pi=\langle V,A,T,s_{0},G\rangle$, where: $V$ is a set of variables with either finite or numeric domains, $A$ is a set of symbolic actions, $T$ is a deterministic transition function, $s_{0}\in S$ is the initial state, and $G$ is the goal description. A state $s\in S$ is a full assignment over the variables in $V$. Since the state space $S$ is typically exponential in $|V|$ (or even infinite in the numeric case), it is not explicitly represented in the input. The goal description $G$ is usually represented as a set of conditions. A state $s_{\ast}\in S$ is considered a goal state if it satisfies $G$, i.e., for each $\varepsilon\in G$ the state $s_{\ast}$ satisfies $\varepsilon$. The transition function $T$ maps a state-action pair to the resulting state, i.e., $T(s,a)=s^{\prime}$. A solution (or plan) for $\Pi$ is a sequence of actions $\langle a_{1},\ldots,a_{n}\rangle$ such that applying them in order leads from $s_{0}$ to a goal state $s_{n}$.
Planning problems can be classified as either satisficing, where the goal is to find any valid plan, or optimal, where the goal is to find a cost-minimizing plan. In this work, we focus on satisficing planning, which prioritizes informative heuristics over theoretically sound ones (e.g., admissibility).
Solutions to planning problems are typically found via heuristic search algorithms such as Greedy Best-First Search (GBFS) or A* (Hart, Nilsson, and Raphael 1968; Pearl 1984). These algorithms are guided by heuristic functions that estimate the cost or distance from a given state to the goal (Pearl 1984). A heuristic function $h:S\to\mathbb{R}_{\geq 0}$ provides such estimates, often derived using domain-independent approximations, such as relaxations of the planning problem. Traditional heuristics are based on symbolic representations and formal action models (e.g., in pddl2.1(Fox and Long 2003), but recent work also explores learned heuristics (Oswald et al. 2024; Chen, Thiébaux, and Trevizan 2024). Informative heuristics greatly reduce the number of explored states, making them essential for practical planning. In this work, we use an LLM to generate a heuristic function based on the formal problem description written in a general-purpose programming language (Rust). This heuristic is integrated into a GBFS planner to obtain satisficing solutions.
Problem Input: pddl2.1 Input: General Description Translate pddl to Rust Generate Rust Functions Generate Heuristic Code via LLM Compile into GBFS and Run Verify with val Verify with Custom Procedure Solution
Figure 1: Procedure flowchart. A general problem description is manually written using Rust components—successor generator, goal test, and initial state—which are then fed into the system. For problems already described in pddl2.1, the Rust translation is derived directly from the encoding.
Methodology
We start with the overview of our approach. The process begins by accepting two types of input: structured representations provided in pddl2.1 (Fox and Long 2003) format and general problem descriptions given informally or in non-standardized formats. The problems specified in pddl2.1 are manually translated into Rust functions that implement a successor generator and a goal-testing function (Izhaki, Green, and Shleyfman 2025). For general problem descriptions, we encode them into Rust according to their specifications. Moreover, the initial and the goal states are specified separately in a JSON file. The manual translation is necessary because current LLMs are not sufficiently reliable for this task, and existing tools lack the maturity required to perform it effectively.
With the foundational components in place, we employ an LLM to generate the heuristic function. The LLM is provided with the Rust code for the successor generator and goal-testing function, along with a prompt designed to produce a heuristic function in Rust suited to the problem. We utilize three prompts, run in succession, for this process. The strategizing prompt [^1] The exact prompt is provided in the supplementary materials. takes as input the domain definition in Rust and outputs a heuristic concept in English. The second prompt, the unrefined heuristic prompt, produces a Rust heuristic reusable across any problem within the domain. Finally, the optional refined heuristic prompt incorporates both the Rust domain file and a JSON file containing specific input, producing a heuristic for the specific problem instance.
The heuristic, successor generator, and goal-testing functions are then integrated into a standard GBFS framework, where the search is guided by the generated heuristic. To ensure correctness, all solutions are verified using appropriate mechanisms based on the input type. For problems specified in pddl2.1, we rely on a standard pddl verifier—val (Howey, Long, and Fox 2004), which cross-checks the solution against the formal specification of the problem. For general problem descriptions, we verify by applying the actions in the plan sequentially, verify the preconditions are met, and that the last state fulfills goal-test. For the visualization of the workflow see Fig. 1. This workflow seamlessly integrates LLM capabilities with automated code generation and classical search techniques, enabling efficient and flexible problem-solving for a wide range of planning domains. By automating key steps, we streamline the planning process and enable the handling of a diverse range of problems with minimal human intervention. The only element that still requires human intervention is the specification of the problem.
Prompt Construction
With the foundational components in place, we employ an LLM to generate the heuristic function. The LLM is provided with the Rust code for the successor generator and goal-testing function. We generate the heuristic using a Chain-of-Thought (Wei et al. 2022) method consisting of three phases (for the full prompts we refer to the Appendix):
Phase 1 (strategizing) [optional]: In the first phase, we give the LLM a Rust domain implementation (successor generator and goal test functions) showing how the domain is simulated and request an English description of a heuristic that could solve it effectively. We further request the model ignore strict theoretical properties and instead will focus on making the heuristic best at facilitating search when combined with GBFS, by approximating monotonous decrease as the given state gets closer to the goal.
Phase 2 (unrefined heuristic): In the second phase, we show precisely the required signature for the heuristic, and request it to implement the english heuristic above into this format. We also provide a list of guidelines to avoid compilation errors, the LLM from requesting further specification, asking us to complete the heuristic, and to deal with other pitfalls. By the end of this phase we have a heuristic which can be slotted into the search.
Phase 3 (refined heuristic) [optional]: In the last phase, we provide the problem instance JSON file, and request a refinement of the heuristic given the details and parameters of the instance. We let the LLM first strategize how it would change the heuristic and only then implement. This allows for more advanced modifications than if it were to immediately start coding autoregressively. We reckon this has four advantages over using the unrefined heuristic:
  1. 1.
    Giving the model another iteration allows it to fix any errors in the initial heuristic.
  2. 2.
    It tells the model how significant each part of the problem is. Consider a simple instance of Pacman without ghosts. If a heuristic includes a complex subroutine to quantify the risk from the ghosts, now it can remove that subroutine. Furthermore, if the subroutine initialized GhostDistance=float::max and iterated over ghosts to find the nearest ghost, the distance may have remained infinity and caused errors, now cleared up by removing the subroutine.
  3. 3.
    The heuristic often contains multiple penalizing elements which are combined with guessed coefficients. Knowing the details of the instance gives the model a better idea on how to tune these coefficients.
  4. 4.
    At times, a solution state is obvious from the initial state. In these cases showing the model the instance details allows it to rewrite the heuristic to guide efficiently toward that solution rather than abstractly guiding toward all solutions. An example of this is given in the supplementary material.
We analyze the impact of including or omitting optional phases in the heuristic generation pipeline in the Ablation Study section.
Baseline Domain-Independent Planners FirstCompilation (fc) TimeSlicedRestarts5 (tsr5)
Rust ENHSP-20 NFD MFF GPT Claude GPT Claude
Domain BFS $h^{\text{md}}$ $^{\text{md}} h+h$$^{\text{mrp}}_{\text{+hj}}$ $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ $h_{\text{LMC}}$ $h_{\text{FF}}$ 4.1 4.1-mini 4o 4o-mini sonnet 3.5 haiku 3.5 4.1 4.1-mini 4o 4o-mini sonnet 3.5 haiku 3.5
Block Grouping (20) 0 9 15 19 18 0 2 13 10 5 2 11 4 18 17 15 2 17 13
Counters (20) 4 11 13 20 11 12 15 10 9 10 9 12 6 11 11 11 12 12 11
Delivery (20) 1 11 17 10 12 9 20 14 2 10 3 9 3 18 15 17 9 20 6
Drone (20) 3 16 20 18 15 16 16 15 12 12 7 12 13 16 17 15 13 17 15
Expedition (20) 3 5 3 0 6 5 5 2 1 1 3 3 1 2 2 3 4 4 2
Farming (20) 3 20 20 20 20 15 9 20 18 17 9 14 18 20 20 20 16 20 20
FO-Counters (20) 4 7 9 0 15 6 20 7 6 4 5 8 4 7 7 7 7 8 7
FO-Farming (20) 5 20 20 15 20 11 16 20 15 19 18 20 15 20 18 20 20 20 20
FO-Sailing (20) 0 0 0 0 1 16 11 8 3 0 0 4 1 17 10 1 0 13 1
Hydropower (20) 20 20 4 0 20 6 1 17 19 19 18 20 11 20 19 20 20 20 20
Market Trader (20) 0 18 17 0 20 0 0 9 3 8 9 1 5 18 9 20 17 7 16
Pathways (20) 0 1 0 13 2 2 13 1 1 1 1 0 0 1 1 1 1 1 1
Plant Watering (20) 1 1 20 20 19 20 13 8 2 2 2 3 1 14 8 10 4 9 4
Rover (20) 2 4 10 8 6 4 10 4 0 2 1 3 0 5 1 6 2 5 3
Sailing (20) 0 0 0 20 17 10 2 9 4 1 0 1 1 18 15 1 0 3 3
Settlers (20) 0 0 0 5 0 0 6 1 1 0 0 0 0 2 1 1 0 1 0
TPP (20) 2 20 20 4 4 2 4 16 3 6 6 11 6 20 12 11 13 20 12
Zenotravel (20) 4 12 14 16 11 10 0 17 11 7 8 11 4 20 18 14 15 17 11
$\sum$ (360) 52 175 202 188 217 144 163 191 120 124 101 143 93 247 201 193 155 214 165
Twin Prime (20) 4 4 - - - - - 16 16 12 7 12 8 19 18 19 14 16 15
Pacman (20) 1 14 - - - - - 14 4 8 2 14 7 16 12 15 12 16 12
$\sum$ (400) 57 193 - - - - - 221 140 144 110 169 108 282 231 227 181 246 192
Table 1: Coverage results of the baseline, domain-independent planners, and the LLM-generated heuristics. We test six LLMs: GPT-4.1, GPT-4.1-mini, GPT-4o, GPT-4o-mini, Claude Sonnet 3.5 and Claude Haiku 3.5. FirstCompilation (fc) and TimeSlicedRestarts5 (tsr5) denote the first compiled heuristic and the five generated heuristics where each heuristic got 100 sec for the run, respectively. All configuration except BFS and MFF (that uses EHCS) use GBFS.
[…]
Figure 2: Per-instance comparison of the runtime of $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ vs. tsr5 GPT-4.1 unrefined. Points below the diagonal favor our approach. On the left we see time spent on solving the instance overall, including the API calls. On the right we display search time only.
Empirical Evaluation
The experiments were conducted on a 13th Gen Intel® Core™ i9-13900 processor running at 2.00 GHz, supported by a 64-bit operating system and 32 GB of RAM. To ensure consistency and reliability, each experiment was constrained by a time limit of 10 minutes and a memory usage limit of 8 GB. We test our approach in the context of satisficing numeric planning. We have evaluated our approach on the domains from the Numeric International Planning Competition (IPC) 2023 (Taitler et al. 2024). [^2] https://github.com/ipc2023-numeric Note that even problems with three unbounded numeric have infinite size state spaces and in general are undecidable (Gnad et al. 2023). As a proof-of-concept, we also include two domains that cannot be easily expressed using pddl:
Twin Prime: The initial state comprises a set of integers and buffers, each holding a single integer. The actions include addition, subtraction, multiplication, and integer division between two registers, with results stored in one register. The goal is to produce a twin prime number exceeding a specified threshold in one of the buffers—a goal that is not easily expressible in pddl.
Deterministic Pacman: The game is defined on a grid containing walls, pellets, ghosts, power-ups, and a single Pacman agent. Pacman moves in one of four cardinal directions per turn, consuming pellets and power-ups upon entry into their cell. The goal is to clear all pellets while avoiding collisions with ghosts that behave in a deterministic manner. Consuming a power-up gives Pacman a temporary ability to banish ghosts and avoid harm from them. This domain cannot be easily expressed in pddl or its extensions due to the complex interactions between Pacman, ghosts, and power-ups, which require temporal dynamics and conditional effects beyond pddl’s representational capabilities.
We evaluated our approach against the following SOTA methods on the IPC domains. All of the planners below support linear expressions in both conditions and assignment effects. As a baseline, we implement a “blind” BFS and GBFS with a manually implemented $h^{\text{md}}$ heuristic (Chen and Thiébaux 2024), running on the Rust implementation of the problems (Izhaki, Green, and Shleyfman 2025). [^3] https://github.com/AlexTuisov/Rust_search
The $h^{\text{md}}$ heuristic estimates the total distance from a state $s$ to satisfying all goal conditions, inspired by the $h^{\text{md}}$ norm. It is defined as:
$h^{\text{md}}(s):=\sum_{\varepsilon\in G}d(\varepsilon,s),$
where $G$ is the set of goal conditions, and $d(\varepsilon,s)$ measures the distance from $s$ to satisfying $\varepsilon$.
[…]
$d(\varepsilon,s):=\inf_{y:\psi(y)\,\bowtie\,0}|\psi(y)-\psi(s)|,$
i.e., the distance from $\psi(s)$ to the nearest satisfying value of $\psi$. Chen and Thiébaux (2024) proposed this approach for linear constraints in pddl2.1, but this heuristic extends naturally to non-linear formulas, though computing $d$ becomes more expensive.
ENHSP-20 is a Java-based planner whose satisficing configurations use GBFS with $h^{\text{mrp}}_{\text{+hj}}$ (Scala et al. 2020), $h^{\text{md}}$ and $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ (Chen and Thiébaux 2024). We used the modified version released by Chen and Thiébaux (2024). [^4] https://github.com/hstairs/jpddlplus/tree/socs24-width-and-mq
Metric-FF (MFF) (Hoffmann 2003) is used off-the-shelf. Implemented in C, it employs Enforced Hill Climbing Search (EHCS) with the interval-relaxed $h_{\text{FF}}$ heuristic and Helpful Operators. [^5] https://fai.cs.uni-saarland.de/hoffmann/metric-ff.html The planner is incomplete and may report problems as unsolvable due to limitations of EHCS.
Numeric-FD (NFD) (Aldinger and Nebel 2017) is a C++ planner that employs Lazy GBFS with numeric LM-cut heuristics (Kuroiwa, Shleyfman, and Beck 2022). [^6] https://github.com/ipc2023-numeric/team-1 This planner won the numeric track of the IPC-2023 (see Taitler et al. 2024).
The Pacman and Twin Prime domains demand a level of expressiveness that makes encoding them for existing planners highly impractical and cumbersome in practice. As there currently are no available methods to derive a domain-independent heuristic for this type of tasks, we benchmarked them against BFS and $h^{\text{md}}$.
LLM-generated heuristics are not always compilable or computationally efficient, occasionally causing the search to exceed memory or time constraints (10 minutes and 8 GB of memory). To address this, we employ two fallback strategies. In FirstCompilation (fc), we repeatedly query the LLM until a compilable heuristic is obtained and run GBFS with it until a solution is found or resources are exhausted, 600 sec (search + API calls). In TimeSlicedRestarts5 (tsr5), we allocate fixed 100 sec time slices to successive heuristics; if a run fails or exceeds memory, we restart with a new heuristic. This continues until a solution is found or the time limit is reached. All API call durations, typically 24–27 seconds, are included in the time budget.
We evaluated four LLMs for this task: OpenAI’s GPT-4.1 and GPT-4.1-mini, and Anthropic’s Claude Sonnet 3.5 and Claude Haiku 3.5. Results are presented in Table 1. As is standard in AI planning, we report coverage rather than averages, since state space size typically grows exponentially and averaging would be misleading. The head-to-head time for solution comparison of tsr5 with GPT-4.1 vs. the best performing $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ can be seen in Figure 2.
Ablation Study
Qry 1 2 3 4 5 6 7 8
4.1
U 199 265 281 281 282 282 282 282
R 162 215 247 258 265 - - -
B - 239 - 275 - 283 - 284
mini
U 194 203 207 228 228 247 247 247
R 121 170 194 210 224 - - -
B - 156 - 205 - 226 - 238
Table 2: Ablation: For GPT-4.1 and GPT-4.1-mini, we compare three configurations: using additional unrefined heuristics only (U), using refined heuristics only (R), and combining both unrefined and refined heuristics (B). The configuration with refined heuristics 6–8 was not run due to time constraints (see Supplementary Materials). We also note that the unrefined version did not improved for up to 10 iterations.
We conducted an ablation study to evaluate the contribution of two components in our heuristic generation pipeline: (i) the heuristic refinement phase and (ii) the preceding strategizing step. The study aimed to determine their impact on heuristic quality and planning performance, as well as the variance in our best configuration. Due to space constraints, results are reported for GPT-4.1; additional configurations are provided in the Supplementary Materials.
Impact of Refinement
The third phase of the pipeline refines the heuristic using instance-specific information from a JSON file. This is achieved by prompting the LLM with both the unrefined heuristic and problem-specific data. We compare the refined heuristics against their unrefined versions.
Empirically, refinement leads to a slight increase in overall coverage (Table 2). However, closer inspection reveals that this improvement stems primarily from increased variance across runs: by combining heuristics generated at different phases (unrefined and refined), the likelihood of obtaining at least one successful variant increases. In other words, the benefit is less a result of systematic improvement and more a consequence of broader sampling.
Impact of Strategizing
We performed an optional first “strategizing” phase, where the LLM is prompted to verbally reason about how to create an effective heuristic for the domain. Reasons for inclusion of this step are described in Methodology section. Surprisingly, in practice we find that despite it’s Chain of Thought step-by-step nature, it actually hurts performance (Table 3).
Letting the model generate code from the start yielded better coverage and required fewer tokens. We observe no consistent improvements in heuristic structure or robustness from allowing strategizing. We attribute this partly to it’s performance on Sailing where it solved 19 instances.
Strat. No Strat.
IPC-23 domains only 247 253
Total 282 288
Avg. input tokens 56.0k 41.3k
Avg. output tokens 8.8k 6.8k
Table 3: Comparison of strategizing vs. no strategizing in terms of token usage and instance counts for TSR5 with GPT-4.1. The average tokens are reported per instance.
Average GPT Claude
TSR 4.1 4.1-mini 4o 4o-mini Sonnet 3.5 Haiku 3.5
Input tokens 56074 71202 66124 72407 76064 90115
Completion tokens 8880 11311 7055 7163 8105 8971
Compilation failure 0.107 0.368 0.122 0.339 0.089 0.373
OutOfMemory rate 0.304 0.241 0.471 0.369 0.448 0.312
Table 4: Token usage and compilation failure rates. We report averages for input tokens and completion tokens per heuristic, and the percentage of compilation failures.
Figure 3: Variance per domain for the tsr5 GPT-4.1 unrefined configuration over three independent runs, each limited to 600 seconds and 8 GB of memory per instance. Total coverages were 282, 282, and 276, indicating stable performance. All runs outperform the domain-independent SOTA. For comparison, we also include results from a run without the strategizing phase.
Summary
The ablation results suggest that refinement has minor positive effects, primarily due to increasing the diversity of heuristic candidates. The strategizing step, however, does not yield measurable benefit. These findings support the use of simple multi-sampling strategies to increase coverage, while questioning the need for more elaborate prompting constructs in heuristic generation.
Discussion
As shown in Table 1, LLM-based heuristics achieve state-of-the-art performance across the evaluated domains. They are particularly effective in challenging benchmarks such as Sailing and Zenotravel, which are known to be difficult for domain-independent planners. The best-performing LLM-based heuristics were produced by GPT-4.1, with Sonnet 3.5 yielding comparable results. Nonetheless, certain domains remain challenging for LLMs, likely due to their sensitivity to non-descriptive variable names and opaque goal specifications, as seen in Pathways and Settlers.
Expressiveness
In the Pacman and Twin Prime domains, no available planner can solve problems requiring this level of expressive power, necessitating a comparison against BFS and $h^{\text{md}}$. As expected, LLM-based heuristics outperformed both. Overall, LLM-based heuristics surpass state-of-the-art numerical planners while addressing problems that cannot be adequately expressed in any existing pddl dialect.
Representation
One concern is that the performance gains observed with LLM-generated heuristics may result from optimized planner code rather than the heuristics themselves. To test this, we reimplemented $h^{\text{md}}$ within our planner for a direct comparison with ENHSP-20 using the same heuristic. As shown in Table 1, ENHSP-20 with $h^{\text{md}}$ generally achieves better results, particularly in domains such as Hydropower and Plant Watering, which are highly sensitive to problem representation. This suggests that departing from pddl-based input to a direct successor-generator representation may, in some cases, degrade performance, allowing us to reject the hypothesis that our representation alone accounts for the observed improvements.
Costs and Efficiency While concerns about the resource efficiency of LLMs have been raised (Stojkovic et al. 2024), energy efficiency is beyond the scope of this work. Nonetheless, our method is amenable to efficient deployment: the first two phases are domain-specific rather than instance-specific and can be executed once per domain, substantially reducing energy and cost. On average, solving an input using TSR5 with GPT-4.1 requires 56k input tokens and 8.8k output tokens (Table 4), corresponding to a pre-caching cost of approximately 0.18 USD.
Conclusion
This paper explored the use of LLMs to generate heuristic functions directly from AI planning task definitions, bypassing traditional domain-independent heuristics written in formal languages like pddl. Our approach uses general-purpose code (Rust) to represent planning tasks and lets LLMs synthesize tailored heuristics.
Empirically, the results show that LLM-generated heuristics achieve state-of-the-art performance across several established planning benchmarks. The approach excels in domains traditionally difficult for domain-independent planners, such as Zenotravel, indicating substantial potential for efficiency gains. Furthermore, it successfully addresses complex planning tasks that are not easily expressible in conventional formalisms, like Pacman and Twin Prime.
Nonetheless, limitations remain. LLM-generated heuristics are sensitive to the clarity and descriptiveness of task representations; unintuitive state-space encodings degrade their quality. Moreover, while larger models such as GPT-4.1 and Sonnet 3.5 tend to produce high-quality heuristics, they incur greater computational costs. Smaller models offer improved cost-efficiency but often generate less reliable code. This trade-off highlights the importance of selecting models according to task complexity and scalability requirements.
Looking forward, this work opens several promising avenues for future research. These include improving robustness across domains, comparing LLM-generated heuristics to hand-crafted ones, and exploring hybrid strategies combining LLMs with domain knowledge or iterative refinement techniques.

### v4

@@
-LLM-Generated Heuristics for AI Planning: Do We Even Need Domain-Independence Anymore?
+Successor-Generator Planning with LLM-generated Heuristics
 Abstract
-Heuristics are a central component of classical AI planning, particularly in domain-independent settings where general applicability is prioritized over task-specific tuning.
-This work revisits that paradigm in light of recent advances in large language models (LLMs), which enable the automatic synthesis of heuristics directly from problem definitions—bypassing the need for handcrafted domain knowledge.
-We present a method that employs LLMs to generate instance-specific heuristic functions from planning tasks specified through successor generators, goal tests, and initial states written in a general-purpose programming language.
+Heuristics are a central component of deterministic planning, particularly in domain-independent settings where general applicability is prioritized over task-specific tuning.
+This work revisits that paradigm in light of recent advances in large language models (LLMs), which enable the automatic synthesis of heuristics directly from problem definitions – bypassing the need for handcrafted domain knowledge.
+We present a method that employs LLMs to generate problem-specific heuristic functions from planning tasks specified through successor generators, goal tests, and initial states written in a general-purpose programming language.
 These heuristics are compiled and integrated into standard heuristic search algorithms, such as greedy best-first search.
@@
 Moreover, it enables the solution of problems that are difficult to express in traditional formalisms, including those with complex numeric constraints or custom transition dynamics.
-We analyze the trade-offs between generality and performance, and argue that LLM-generated heuristics represent a promising direction for integrating machine learning with symbolic reasoning in planning.
+We provide an extensive empirical evaluation that characterizes the strengths and limitations of the approach across diverse planning settings, demonstrating its effectiveness.
 Introduction
-Deterministic AI planning, at its core, involves finding a sequence of actions that transitions an agent from an initial state to a goal state, given a formal description of the environment’s dynamics.
-To search effectively through large state spaces, planners rely on heuristics—functions that estimate the distance to the goal and guide the search toward promising regions.
-Classical planning has traditionally focused on general, domain- and problem-independent heuristics, which offer broad applicability and avoid the need for task-specific design (Bonet and Geffner 2001).
-However, the performance of such heuristics can depend on the representation language and the structural characteristics of the domain.
-In more expressive settings, general heuristics may face limitations or require adaptation to remain effective in different versions of pddl (cf.
-Fox and Long 2003; Edelkamp and Hoffmann 2004; Gerevini and Long 2005).
-In this work, we propose to automatically generate task-specific heuristic functions from formal problem definitions, using LLMs to generate task-specific heuristics, that are then used by a classic AI search algorithm.
-In this way, we benefit from the systematic and guaranteed search properties of the classic algorithm, while keeping the generation of the heuristics automatic, with no need for expert manual labor for heuristics design.
-We represent each planning task using three components implemented in a general-purpose programming language: a successor generator, a goal test, and an initial state.
-These are provided as input to the LLM, which returns a heuristic function which can be tailored to the specific problem structure.
-This heuristic is compiled and used to guide search without further interaction with the LLM.
-This approach has several advantages.
-First, it avoids repeated LLM calls during planning, greatly improving efficiency.
-Second, it enables the use of transparent, verifiable search components: both the LLM-generated heuristic and the underlying task specification are standard Rust code that can be debugged and tested.
-Third, it allows us to solve planning problems that involve constructs difficult to express in traditional frameworks—such as recursive goals, complex numeric conditions, or custom transition logic.
-We show that LLMs can be used to improve general-purpose planners by using them to generate input-specific heuristics, rather than relying on general-purpose heuristics.
-We evaluate our approach on a set of standard AI planning benchmarks, translated into Rust representations suitable for LLM prompting.
-In many domains, the resulting heuristics outperform classical domain-independent methods and compete with state-of-the-art numeric planners.
-Additionally, our method is able to solve novel planning tasks that lack formal encoding in existing planning languages.
+Deterministic AI planning involves finding a sequence of actions that moves an agent from an initial state to a goal state, given a formal description of the environment’s dynamics (ghallab-et-al:2004).
+Because state spaces are usually exponential or even infinite, effective search strategies require heuristics – functions that estimate the distance to the goal and help prioritize promising states (pearl:1984; bonet2001planning).
+Heuristic search has therefore become a central paradigm in classical planning, underpinning many of the most successful planners developed over the past decades (hoffman:jair-2003; helmert:jair-2006; scala-et-al:ijcai-2016; aldinger-nebel:socs-2017).
+Classical planning has traditionally focused on domain-independent heuristics, which offer broad applicability and avoid the need for task-specific design (bonet2001planning).
+However, the performance of such heuristics can vary with the representation language and the structural characteristics of the domain.
+In more expressive settings, general heuristics may encounter challenges and need adaptation to remain effective across different versions of pddl (mcdermott-et-al:tr-1998; fox-long:jair-2003; edelkamp-hoffmann:tr-2004) (e.g., cf.
+classical vs.
+numeric settings (hoffman:aimag-2001; hoffman:jair-2003; helmert-domshlak:icaps-2009; kuroiwa-et-al:jair-2022)).
+For further discussion of current limitations within pddl see edelkamp:icapsws-2003; rintanen:aaai-20215.
+In this work, we propose a method for automatically generating problem-specific heuristic functions from formal problem definitions using LLMs.
+The generated heuristics are then used by a sound search algorithm, such as greedy best-first search (GBFS).
+This setup preserves the systematic and provably correct behavior of classical search methods, while making heuristic construction fully automatic and eliminating the need for expert-designed heuristics.
+We represent each planning task using three components, implemented in a general-purpose programming language: a successor generator, a goal test, and an initial state (russell-norvig:1995; guan-etal:nips-2023; oswald-etal:icaps-2024).
+We refer to this representation as an Explicit Successor Generator (ESG).
+These ESG components, together with a prompt, are provided as input to the LLM, which returns a heuristic function (expressed in the same programming language) tailored to the structure of the given problem.
+The heuristic is then compiled and used to guide search without any further interaction with the LLM.
+This approach offers several benefits.
+It removes the need for repeated LLM calls during planning—a source of significant inefficiency in prior work (tos:nips-2024; valmeekam-etal:corr-2022; kambhampati-etal:icml-2024).
+It also provides transparent search components: both the LLM-generated heuristic and the task specification are expressed in a general-purpose programming language (Rust in this work), allowing inspection, debugging, and testing.
+In addition, it accommodates planning tasks that involve constructs difficult to model in traditional frameworks, including recursive goals, intricate numeric conditions, the creation of new variables, or custom transition logic.
+We show that LLM-generated heuristic can strengthen general-purpose ESG planning.
+Our evaluation includes a set of standard numerical planning benchmarks represented in the ESG form (using Rust), as well as new benchmarks that are challenging to encode in pddl.
+Across many domains, the resulting heuristics achieve strong performance, matching the level of state-of-the-art numeric planners in numerous cases.
+Furthermore, the method is able to solve planning tasks that currently lack formal encodings in existing planning languages.
 Related Work
-This paper revisits AI planning by heuristic search, by replacing task-independent heuristics with LLM-generated task-specific heuristics.
+This paper revisits AI planning by heuristic search, by replacing domain-independent heuristics with LLM-generated domain-specific ones.
 Below, we review related work on LLMs and planning.
-For a more comprehensive overview of LLMs as planning modelers, see the surveys by (Tantakoun, Zhu, and Muise 2025) and by (Aghzal et al.
-2025).
+For a more comprehensive overview of LLMs as planning modelers, see the surveys by (tantakoun-etal:corr-2025) and by (aghzal2025surveylargelanguagemodels).
 Planning with LLMs
-(Valmeekam et al.
-2022) provides an important early evaluation of LLM-based planning, demonstrating that even on simple benchmarks, LLMs do not consistently exhibit reliable planning behavior.
+(valmeekam-etal:corr-2022) provides an important early evaluation of LLM-based planning, demonstrating that even on simple benchmarks, LLMs do not consistently exhibit reliable planning behavior.
 While it seems that LLMs cannot yet generate plans efficiently, our work shows that they can be used effectively as a component of a larger planning paradigm.
 Namely, LLMs can generate heuristics and search components rather than perform planning directly.
-A closely related approach is presented by (Katz et al.
-2024), who argue that existing LLM-based planning methods are neither sound nor complete and incur high computational costs due to repeated LLM calls.
+A closely related approach is presented by (tos:nips-2024), who argue that existing LLM-based planning methods are neither sound nor complete and incur high computational costs due to repeated LLM calls.
 They propose an alternative, Thought of Search (ToS), in which LLMs are used to generate symbolic search components—successor functions and goal tests.
@@
 Although their work focuses on small, finite search problems such as the 24 game and Mini Crosswords, they provide compelling evidence that such an approach can drastically reduce computational cost and improve correctness (solved by BFS).
-The work shows that with respect to planning their approach overperforms the planning directly with LLM approaches such as: Chain-of-Thought (Wei et al.
-2022), ReAct (Yao et al.
-2023b), ReWOO (Xu et al.
-2023), Reasoning via Planning (Hao et al.
-2023), Tree of Thoughts (Yao et al.
-2023a), Graph of Thoughts (Besta et al.
-2024), Reflexion (Shinn et al.
-2023), Algorithm of Thought (Sel et al.
-2024).
+The work shows that with respect to planning their approach overperforms the planning directly with LLM approaches such as: Chain-of-Thought (wei2022chain), ReAct (react:iclr-2023), ReWOO (rewoo:corr-2023), Reasoning via Planning (rap:emnkp-2023), Tree of Thoughts (tot:nips-2023), Graph of Thoughts (got:aaai-2024), Reflexion (reflection:nips-2023), Algorithm of Thought (aot:icml-2024).
 The ToS approach runs a breadth-first search (BFS) using components—successor functions and goal tests generated by LLM.
 Our approach extends this intuition by employing heuristic search methods to address problems of significantly greater complexity than those solvable through naive BFS—specifically in AI planning, where transition systems often become infinite due to the presence of unbounded numeric fluents.
-Consequently, generating effective heuristics is particularly crucial, underscoring that the symbolic utilization of LLMs can be successfully applied beyond toy domains.
-A recent study in a similar vein to our work is presented in (Corrêa, Pereira, and Seipp 2025).
-While their approach targets planning tasks specified in the Planning Domain Definition Language (PDDL), ours departs from this framework by operating directly on successor generators and goal-check functions.
+Thus, generating effective heuristics is particularly crucial, underscoring that the symbolic utilization of LLMs can be successfully applied beyond toy domains.
+A recent study in a similar vein to our work is presented in (correa2025classical).
+While their approach targets planning tasks specified in pddl, ours departs from this framework by operating directly on successor generators and goal-check functions.
 This shift affords greater representational flexibility, enabling the modeling of tasks that PDDL cannot express effectively.
@@
 Heuristic Generation Outside of Planning
-LLMs have also been explored for heuristic generation beyond automated planning, notably in domains such as online bin packing and TSP (Liu et al.
-2024), as well as other combinatorial optimization problems (Ye et al.
-2024).
+LLMs have also been explored for heuristic generation beyond automated planning, notably in domains such as online bin packing and TSP (liu2024evolution), as well as other combinatorial optimization problems (ye2024reevo).
 These approaches typically combine LLMs with evolutionary algorithms to evolve heuristics that guide the optimization process.
@@
 Problem Generation
-Recent work has explored using LLMs to generate domains and problem descriptions using pddl2.1 (Fox and Long 2003), among such works are (Guan et al.
-2023; Liu et al.
-2023; Oswald et al.
-2024).
-For instance, (Silver et al.
-2024) demonstrated that GPT-4 can output plans from parsed pddl2.1 inputs using simple, non-search-based strategies.
-(Guan et al.
-2023) proposed an approach in which pddl2.1 actions are generated one at a time via LLM queries, refined using human feedback.
-A related concept appears in (Zhou et al.
-2024), who use LLMs to generate, execute, and refine Python code for solving mathematical reasoning problems.
+Recent work has explored using LLMs to generate domains and problem descriptions using pddl2.1 (fox-long:jair-2003), among such works are (guan-etal:nips-2023; liu-etal:corr-2023; oswald-etal:icaps-2024).
+For instance, (silver-etal:aaai-2024) demonstrated that GPT-4 can output plans from parsed pddl2.1 inputs using simple, non-search-based strategies.
+(guan-etal:nips-2023) proposed an approach in which pddl2.1 actions are generated one at a time via LLM queries, refined using human feedback.
+A related concept appears in (zhou-etal:iclr-2024), who use LLMs to generate, execute, and refine Python code for solving mathematical reasoning problems.
 Their findings support the idea that LLMs can produce verifiable code with minimal manual intervention.
@@
 Code Generation Our approach relies on LLMs not to solve problems directly, but to generate search guidance in the form of state evaluation heuristic code.
-(Zhou et al.
-2024) pursued a similar strategy in mathematical domains, showing that LLMs can generate and iteratively improve executable code.
+(zhou-etal:iclr-2024) pursued a similar strategy in mathematical domains, showing that LLMs can generate and iteratively improve executable code.
 Our work builds on this insight, but targets search guidance functions within a planning framework.
-The viability of this approach is supported by recent advances in program synthesis and LLM-based code generation (Madaan et al.
-2023; Zhang et al.
-2023; Chen et al.
-2024; Muennighoff et al.
-2024; Zhong, Wang, and Shang 2024), which show that language models are increasingly capable of producing correct, type-safe, and semantically meaningful programs.
+The viability of this approach is supported by recent advances in program synthesis and LLM-based code generation (madaan-etal:nips-2023; zhang-etal:acl-2023; chen-etal:iclr-2024; muennighoff-etal:iclr-2024; zhong-etal:aclf-2024), which show that language models are increasingly capable of producing correct, type-safe, and semantically meaningful programs.
 Deterministic AI Planning
@@
 In this work, we focus on satisficing planning, which prioritizes informative heuristics over theoretically sound ones (e.g., admissibility).
-Solutions to planning problems are typically found via heuristic search algorithms such as Greedy Best-First Search (GBFS) or A* (Hart, Nilsson, and Raphael 1968; Pearl 1984).
-These algorithms are guided by heuristic functions that estimate the cost or distance from a given state to the goal (Pearl 1984).
+Solutions to planning problems are typically found via heuristic search algorithms such as GBFS or A* (hart-et-al:ieee-ssc-1968; pearl:1984).
+These algorithms are guided by heuristic functions that estimate the cost or distance from a given state to the goal (pearl:1984).
 A heuristic function $h:S\to\mathbb{R}_{\geq 0}$ provides such estimates, often derived using domain-independent approximations, such as relaxations of the planning problem.
-Traditional heuristics are based on symbolic representations and formal action models (e.g., in pddl2.1(Fox and Long 2003), but recent work also explores learned heuristics (Oswald et al.
-2024; Chen, Thiébaux, and Trevizan 2024).
+Traditional heuristics are based on symbolic representations and formal action models (e.g., in pddl2.1(fox-long:jair-2003), but recent work also explores learned heuristics (oswald-etal:icaps-2024; chen-etal:aaai-2024).
 Informative heuristics greatly reduce the number of explored states, making them essential for practical planning.
@@
 We start with the overview of our approach.
-The process begins by accepting two types of input: structured representations provided in pddl2.1 (Fox and Long 2003) format and general problem descriptions given informally or in non-standardized formats.
-The problems specified in pddl2.1 are manually translated into Rust functions that implement a successor generator and a goal-testing function (Izhaki, Green, and Shleyfman 2025).
+The process begins by accepting two types of input: structured representations provided in pddl2.1 (fox-long:jair-2003) format and general problem descriptions given informally or in non-standardized formats.
+The problems specified in pddl2.1 are manually translated into Rust functions that implement a successor generator and a goal-testing function.
 For general problem descriptions, we encode them into Rust according to their specifications.
@@
 The manual translation is necessary because current LLMs are not sufficiently reliable for this task, and existing tools lack the maturity required to perform it effectively.
+The automatic translations to ESG formalism is left for future work.
 With the foundational components in place, we employ an LLM to generate the heuristic function.
 The LLM is provided with the Rust code for the successor generator and goal-testing function, along with a prompt designed to produce a heuristic function in Rust suited to the problem.
-We utilize three prompts, run in succession, for this process.
-The strategizing prompt [^1] The exact prompt is provided in the supplementary materials.
-takes as input the domain definition in Rust and outputs a heuristic concept in English.
-The second prompt, the unrefined heuristic prompt, produces a Rust heuristic reusable across any problem within the domain.
-Finally, the optional refined heuristic prompt incorporates both the Rust domain file and a JSON file containing specific input, producing a heuristic for the specific problem instance.
+The prompt [^1] The exact prompt for both variations is given in the Appendix.
+has three components:
+  1. 1.
+    Conditioning the model to be a senior engineer in the GPL used (Anam2025PromptEA) and providing the format the resulting heuristic must follow.
+This component is spread across both the system prompt and the user message [^2] We include guidance to avoid compilation errors, however the guidance could be far better optimized.
+This optimization, however, is out of scope for this work.
+  2. 2.
+    Requesting that the LLM generate a heuristic.
+  3. 3.
+    Providing the model with the GPL domain implementation (the ESG).
+As an additional variation of our method, the user message may also contain a fourth component:
+  1. 4.
+    The JSON representation of the instance to be solved.
+We analyze the impact of instance-specificity in the heuristic generation pipeline in the Ablation Study section.
 The heuristic, successor generator, and goal-testing functions are then integrated into a standard GBFS framework, where the search is guided by the generated heuristic.
 To ensure correctness, all solutions are verified using appropriate mechanisms based on the input type.
-For problems specified in pddl2.1, we rely on a standard pddl verifier—val (Howey, Long, and Fox 2004), which cross-checks the solution against the formal specification of the problem.
+For problems specified in pddl2.1, we rely on a standard pddl verifier—val (howey-etal:ictai-2004), which cross-checks the solution against the formal specification of the problem.
 For general problem descriptions, we verify by applying the actions in the plan sequentially, verify the preconditions are met, and that the last state fulfills goal-test.
@@
 1.
-This workflow seamlessly integrates LLM capabilities with automated code generation and classical search techniques, enabling efficient and flexible problem-solving for a wide range of planning domains.
-By automating key steps, we streamline the planning process and enable the handling of a diverse range of problems with minimal human intervention.
-The only element that still requires human intervention is the specification of the problem.
-Prompt Construction
-With the foundational components in place, we employ an LLM to generate the heuristic function.
-The LLM is provided with the Rust code for the successor generator and goal-testing function.
-We generate the heuristic using a Chain-of-Thought (Wei et al.
-2022) method consisting of three phases (for the full prompts we refer to the Appendix):
-Phase 1 (strategizing) [optional]: In the first phase, we give the LLM a Rust domain implementation (successor generator and goal test functions) showing how the domain is simulated and request an English description of a heuristic that could solve it effectively.
-We further request the model ignore strict theoretical properties and instead will focus on making the heuristic best at facilitating search when combined with GBFS, by approximating monotonous decrease as the given state gets closer to the goal.
-Phase 2 (unrefined heuristic): In the second phase, we show precisely the required signature for the heuristic, and request it to implement the english heuristic above into this format.
-We also provide a list of guidelines to avoid compilation errors, the LLM from requesting further specification, asking us to complete the heuristic, and to deal with other pitfalls.
-By the end of this phase we have a heuristic which can be slotted into the search.
-Phase 3 (refined heuristic) [optional]: In the last phase, we provide the problem instance JSON file, and request a refinement of the heuristic given the details and parameters of the instance.
-We let the LLM first strategize how it would change the heuristic and only then implement.
-This allows for more advanced modifications than if it were to immediately start coding autoregressively.
-We reckon this has four advantages over using the unrefined heuristic:
-  1. 1.
-    Giving the model another iteration allows it to fix any errors in the initial heuristic.
-  2. 2.
-    It tells the model how significant each part of the problem is.
-Consider a simple instance of Pacman without ghosts.
-If a heuristic includes a complex subroutine to quantify the risk from the ghosts, now it can remove that subroutine.
-Furthermore, if the subroutine initialized GhostDistance=float::max and iterated over ghosts to find the nearest ghost, the distance may have remained infinity and caused errors, now cleared up by removing the subroutine.
-  3. 3.
-    The heuristic often contains multiple penalizing elements which are combined with guessed coefficients.
-Knowing the details of the instance gives the model a better idea on how to tune these coefficients.
-  4. 4.
-    At times, a solution state is obvious from the initial state.
-In these cases showing the model the instance details allows it to rewrite the heuristic to guide efficiently toward that solution rather than abstractly guiding toward all solutions.
-An example of this is given in the supplementary material.
-We analyze the impact of including or omitting optional phases in the heuristic generation pipeline in the Ablation Study section.
-Baseline Domain-Independent Planners FirstCompilation (fc) TimeSlicedRestarts5 (tsr5)
-Rust ENHSP-20 NFD MFF GPT Claude GPT Claude
-Domain BFS $h^{\text{md}}$ $^{\text{md}} h+h$$^{\text{mrp}}_{\text{+hj}}$ $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ $h_{\text{LMC}}$ $h_{\text{FF}}$ 4.1 4.1-mini 4o 4o-mini sonnet 3.5 haiku 3.5 4.1 4.1-mini 4o 4o-mini sonnet 3.5 haiku 3.5
-Block Grouping (20) 0 9 15 19 18 0 2 13 10 5 2 11 4 18 17 15 2 17 13
-Counters (20) 4 11 13 20 11 12 15 10 9 10 9 12 6 11 11 11 12 12 11
-Delivery (20) 1 11 17 10 12 9 20 14 2 10 3 9 3 18 15 17 9 20 6
-Drone (20) 3 16 20 18 15 16 16 15 12 12 7 12 13 16 17 15 13 17 15
-Expedition (20) 3 5 3 0 6 5 5 2 1 1 3 3 1 2 2 3 4 4 2
-Farming (20) 3 20 20 20 20 15 9 20 18 17 9 14 18 20 20 20 16 20 20
-FO-Counters (20) 4 7 9 0 15 6 20 7 6 4 5 8 4 7 7 7 7 8 7
-FO-Farming (20) 5 20 20 15 20 11 16 20 15 19 18 20 15 20 18 20 20 20 20
-FO-Sailing (20) 0 0 0 0 1 16 11 8 3 0 0 4 1 17 10 1 0 13 1
-Hydropower (20) 20 20 4 0 20 6 1 17 19 19 18 20 11 20 19 20 20 20 20
-Market Trader (20) 0 18 17 0 20 0 0 9 3 8 9 1 5 18 9 20 17 7 16
-Pathways (20) 0 1 0 13 2 2 13 1 1 1 1 0 0 1 1 1 1 1 1
-Plant Watering (20) 1 1 20 20 19 20 13 8 2 2 2 3 1 14 8 10 4 9 4
-Rover (20) 2 4 10 8 6 4 10 4 0 2 1 3 0 5 1 6 2 5 3
-Sailing (20) 0 0 0 20 17 10 2 9 4 1 0 1 1 18 15 1 0 3 3
-Settlers (20) 0 0 0 5 0 0 6 1 1 0 0 0 0 2 1 1 0 1 0
-TPP (20) 2 20 20 4 4 2 4 16 3 6 6 11 6 20 12 11 13 20 12
-Zenotravel (20) 4 12 14 16 11 10 0 17 11 7 8 11 4 20 18 14 15 17 11
-$\sum$ (360) 52 175 202 188 217 144 163 191 120 124 101 143 93 247 201 193 155 214 165
-Twin Prime (20) 4 4 - - - - - 16 16 12 7 12 8 19 18 19 14 16 15
-Pacman (20) 1 14 - - - - - 14 4 8 2 14 7 16 12 15 12 16 12
-$\sum$ (400) 57 193 - - - - - 221 140 144 110 169 108 282 231 227 181 246 192
-Table 1: Coverage results of the baseline, domain-independent planners, and the LLM-generated heuristics.
-We test six LLMs: GPT-4.1, GPT-4.1-mini, GPT-4o, GPT-4o-mini, Claude Sonnet 3.5 and Claude Haiku 3.5.
-FirstCompilation (fc) and TimeSlicedRestarts5 (tsr5) denote the first compiled heuristic and the five generated heuristics where each heuristic got 100 sec for the run, respectively.
-All configuration except BFS and MFF (that uses EHCS) use GBFS.
+This workflow integrates LLM capabilities with automated code generation and classical search techniques, enabling efficient and flexible problem-solving for a wide range of planning domains.
+Note that the only element of the workflow that still requires human intervention is the specification of the problem.
+Baseline Domain-Independent Planners FirstCompilation (fc) UntilSuccess (US) SelfPortfolio-10 (SP-10)
+Rust ENHSP-20 NFD MFF GPT Claude GPT Claude GPT Claude
+Domain BFS $h^{\text{md}}$ $h^{\text{md}}$ $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ $P(3h||3n)^{\ddagger}$ $h_{\text{LMC}}$ $h_{\text{FF}}$ 4.1 5.1 Sonnet 4.5 4.1 5.1 Sonnet 4.5 4.1 5.1 Sonnet 4.5
 […]
-Figure 2: Per-instance comparison of the runtime of $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ vs.
-tsr5 GPT-4.1 unrefined.
-Points below the diagonal favor our approach.
-On the left we see time spent on solving the instance overall, including the API calls.
-On the right we display search time only.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 1: Coverage results of the baseline, domain-independent planners, the best portfolio, and the LLM-generated heuristics.
+We report our results from GPT-4.1, GPT-5.1, and Claude Sonnet 4.5 with configurations FirstCompilation (fc), UntilSuccess (us), and SelfPortfolio-10 (SP-10).
+All configuration except BFS and MFF (which uses EHCS) use GBFS.
+Models marked with are reasoning models.
+For results on all models and configurations see Appendix.
+The $P(3h||3n)$ portfolio uniformly divides time among three “best” ENHSP heuristics and three “best” novelty heuristics (chen-thiebaux:socs-2024), where the notion of “best” is derived from performance on the same problem set used in our evaluation, rather than being fixed independently of it.
 Empirical Evaluation
@@
 We test our approach in the context of satisficing numeric planning.
-We have evaluated our approach on the domains from the Numeric International Planning Competition (IPC) 2023 (Taitler et al.
-2024).
-[^2] https://github.com/ipc2023-numeric Note that even problems with three unbounded numeric have infinite size state spaces and in general are undecidable (Gnad et al.
-2023).
+We have evaluated our approach on the domains from the Numeric International Planning Competition (IPC) 2023 (taitler-etal:aimag-2024).
+[^3] https://github.com/ipc2023-numeric Note that even problems with three unbounded numeric have infinite size state spaces and in general are undecidable (gnad-et-al:icaps-2023).
 As a proof-of-concept, we also include two domains that cannot be easily expressed using pddl:
@@
 This domain cannot be easily expressed in pddl or its extensions due to the complex interactions between Pacman, ghosts, and power-ups, which require temporal dynamics and conditional effects beyond pddl’s representational capabilities.
+(a) Total Time (including failed generation + run pairs)
+(b) Expanded States
+Figure 2: Per-instance comparisons of the Total Time (left) and expanded states (right) between GPT-4.1 with SelfPortfolio-10 and $h^{\text{add}}_{\langle\text{B, QB}\rangle}$.
+Points below the diagonal favor our approach.
+On problems both solve they seem to provide similar levels of heuristic guidance.
+[^5] A portfolio between the two achieves a coverage of 304, however this is a post-hoc optimization and not a fair comparison.
+We do however consider Domain-Independent+LLM a fruitful direction fo future work.
+Figure 3: Per-instance comparison of the Total Time (generation + run) (up) and expanded states (down) between GPT-5.1 with and without setting the reasoning_effort to high.
+Points below the diagonal favor high reasoning effort.
+Allowing increased reasoning effort moderately increases heuristic quality, but comes at a heavy expense of time.
 We evaluated our approach against the following SOTA methods on the IPC domains.
 All of the planners below support linear expressions in both conditions and assignment effects.
-As a baseline, we implement a “blind” BFS and GBFS with a manually implemented $h^{\text{md}}$ heuristic (Chen and Thiébaux 2024), running on the Rust implementation of the problems (Izhaki, Green, and Shleyfman 2025).
-[^3] https://github.com/AlexTuisov/Rust_search
-The $h^{\text{md}}$ heuristic estimates the total distance from a state $s$ to satisfying all goal conditions, inspired by the $h^{\text{md}}$ norm. It is defined as:
+As a baseline, we implement a “blind” BFS and GBFS with a manually implemented $h^{\text{md}}$ heuristic (chen-thiebaux:socs-2024), running on a Rust implementation of the problems based on (Domains_Project).
+The $h^{\text{md}}$ heuristic estimates the total distance from a state $s$ to satisfying all goal conditions, inspired by the $h^{\text{md}}$ norm:
 $h^{\text{md}}(s):=\sum_{\varepsilon\in G}d(\varepsilon,s),$
@@
 i.e., the distance from $\psi(s)$ to the nearest satisfying value of $\psi$.
-Chen and Thiébaux (2024) proposed this approach for linear constraints in pddl2.1, but this heuristic extends naturally to non-linear formulas, though computing $d$ becomes more expensive.
-ENHSP-20 is a Java-based planner whose satisficing configurations use GBFS with $h^{\text{mrp}}_{\text{+hj}}$ (Scala et al.
-2020), $h^{\text{md}}$ and $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ (Chen and Thiébaux 2024).
-We used the modified version released by Chen and Thiébaux (2024).
-[^4] https://github.com/hstairs/jpddlplus/tree/socs24-width-and-mq
-Metric-FF (MFF) (Hoffmann 2003) is used off-the-shelf.
-Implemented in C, it employs Enforced Hill Climbing Search (EHCS) with the interval-relaxed $h_{\text{FF}}$ heuristic and Helpful Operators.
-[^5] https://fai.cs.uni-saarland.de/hoffmann/metric-ff.html The planner is incomplete and may report problems as unsolvable due to limitations of EHCS.
-Numeric-FD (NFD) (Aldinger and Nebel 2017) is a C++ planner that employs Lazy GBFS with numeric LM-cut heuristics (Kuroiwa, Shleyfman, and Beck 2022).
-[^6] https://github.com/ipc2023-numeric/team-1 This planner won the numeric track of the IPC-2023 (see Taitler et al.
-2024).
+chen-thiebaux:socs-2024 proposed this approach for linear constraints in pddl2.1, but this heuristic extends naturally to non-linear formulas, though computing $d$ becomes more expensive.
+ENHSP-20 is a Java-based planner whose satisficing configurations use GBFS by (scala-et-al:jair-2020), using the modified version released by chen-thiebaux:socs-2024.666github.com/hstairs/jpddlplus/tree/socs24-width-and-mq.
+Following (chen-thiebaux:socs-2024) we compare against $^{\text{md}}, h+h$$^{\text{mrp}}_{\text{+hj}}$, $h^{\text{add}}$, their novelty variants $^{\text{md}}_{\langle\text{B, QB}\rangle}, h+h$$^{\text{mrp}}_{\text{+hj}~\langle\text{B, QB}\rangle}$, $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, as well as the three portfolio settings, $P(3h)$ (non-novelty only), $P(3n)$ (novelty only), and $P(3h\,\|\,3n)$ (both).
+We report of them only $h^{\text{md}}$, the best-performing single heuristic $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, and the best portfolio configuration, $P(3h\,\|\,3n)$, while the full comparison is provided in Appendix.
+Metric-FF (MFF) (hoffman:jair-2003) is used off-the-shelf.
+Implemented in C, it employs Enforced Hill Climbing Search (EHCS) with the interval-relaxed $h_{\text{FF}}$ heuristic and Helpful Operators.777fai.cs.uni-saarland.de/hoffmann/metric-ff.html The planner is incomplete and may report problems as unsolvable due to limitations of EHCS.
+Numeric-FD (NFD) (aldinger-nebel:socs-2017) is a C++ planner that employs Lazy GBFS with numeric LM-cut heuristics (kuroiwa-et-al:icaps-2022).888github.com/ipc2023-numeric/team-1 This planner won the numeric track of the IPC-2023 (see taitler-etal:aimag-2024).
 The Pacman and Twin Prime domains demand a level of expressiveness that makes encoding them for existing planners highly impractical and cumbersome in practice.
 As there currently are no available methods to derive a domain-independent heuristic for this type of tasks, we benchmarked them against BFS and $h^{\text{md}}$.
-LLM-generated heuristics are not always compilable or computationally efficient, occasionally causing the search to exceed memory or time constraints (10 minutes and 8 GB of memory).
-To address this, we employ two fallback strategies.
-In FirstCompilation (fc), we repeatedly query the LLM until a compilable heuristic is obtained and run GBFS with it until a solution is found or resources are exhausted, 600 sec (search + API calls).
-In TimeSlicedRestarts5 (tsr5), we allocate fixed 100 sec time slices to successive heuristics; if a run fails or exceeds memory, we restart with a new heuristic.
-This continues until a solution is found or the time limit is reached.
-All API call durations, typically 24–27 seconds, are included in the time budget.
-We evaluated four LLMs for this task: OpenAI’s GPT-4.1 and GPT-4.1-mini, and Anthropic’s Claude Sonnet 3.5 and Claude Haiku 3.5.
+LLM-generated heuristics are not always compilable and vary in quality and efficiency generation-to-generation. To address this, we employ three fallback strategies:
+In FirstCompilation (fc), we repeatedly query the LLM until a compilable heuristic is obtained and run GBFS with it until a solution is found or resources are exhausted, 600 sec (search + all API calls).
+In UntilSuccess (us), we operate the same as fc, except we continue even if a compilable heuristic failed.
+In SelfPortfolio-N (sp-n), we allocate fixed $Time/N$s slices to N heuristics; if a run fails or exceeds memory, we restart with generating a new heuristic.
+This continues until a solution is found or all slices failed.
+All API call durations, typically 10–45 seconds depending on the model, are included in the time budget.
+We evaluated The following LLMs and settings: OpenAI’s GPT-4.1 and GPT-4.1-mini, their reasoning models GPT-5.1 and GPT-5-mini (under both a ”low” and ”high” reasoning effort setting), and Anthropic’s Claude Sonnet 4.5 and Claude Haiku 4.5.
 Results are presented in Table 1.
+For practicality, we limit the maximum number of heuristics generated for fc and us to 10, which if passed the instance is declared failed.
 As is standard in AI planning, we report coverage rather than averages, since state space size typically grows exponentially and averaging would be misleading.
-The head-to-head time for solution comparison of tsr5 with GPT-4.1 vs.
-the best performing $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ can be seen in Figure 2.
-Ablation Study
-Qry 1 2 3 4 5 6 7 8
-4.1
-U 199 265 281 281 282 282 282 282
-R 162 215 247 258 265 - - -
-B - 239 - 275 - 283 - 284
-mini
-U 194 203 207 228 228 247 247 247
-R 121 170 194 210 224 - - -
-B - 156 - 205 - 226 - 238
-Table 2: Ablation: For GPT-4.1 and GPT-4.1-mini, we compare three configurations: using additional unrefined heuristics only (U), using refined heuristics only (R), and combining both unrefined and refined heuristics (B).
-The configuration with refined heuristics 6–8 was not run due to time constraints (see Supplementary Materials).
-We also note that the unrefined version did not improved for up to 10 iterations.
-We conducted an ablation study to evaluate the contribution of two components in our heuristic generation pipeline: (i) the heuristic refinement phase and (ii) the preceding strategizing step.
-The study aimed to determine their impact on heuristic quality and planning performance, as well as the variance in our best configuration.
-Due to space constraints, results are reported for GPT-4.1; additional configurations are provided in the Supplementary Materials.
-Impact of Refinement
-The third phase of the pipeline refines the heuristic using instance-specific information from a JSON file.
-This is achieved by prompting the LLM with both the unrefined heuristic and problem-specific data.
-We compare the refined heuristics against their unrefined versions.
-Empirically, refinement leads to a slight increase in overall coverage (Table 2).
-However, closer inspection reveals that this improvement stems primarily from increased variance across runs: by combining heuristics generated at different phases (unrefined and refined), the likelihood of obtaining at least one successful variant increases.
-In other words, the benefit is less a result of systematic improvement and more a consequence of broader sampling.
-Impact of Strategizing
-We performed an optional first “strategizing” phase, where the LLM is prompted to verbally reason about how to create an effective heuristic for the domain.
-Reasons for inclusion of this step are described in Methodology section.
-Surprisingly, in practice we find that despite it’s Chain of Thought step-by-step nature, it actually hurts performance (Table 3).
-Letting the model generate code from the start yielded better coverage and required fewer tokens.
-We observe no consistent improvements in heuristic structure or robustness from allowing strategizing.
-We attribute this partly to it’s performance on Sailing where it solved 19 instances.
-Strat. No Strat.
-IPC-23 domains only 247 253
-Total 282 288
-Avg. input tokens 56.0k 41.3k
-Avg. output tokens 8.8k 6.8k
-Table 3: Comparison of strategizing vs. no strategizing in terms of token usage and instance counts for TSR5 with GPT-4.1. The average tokens are reported per instance.
-Average GPT Claude
-TSR 4.1 4.1-mini 4o 4o-mini Sonnet 3.5 Haiku 3.5
-Input tokens 56074 71202 66124 72407 76064 90115
-Completion tokens 8880 11311 7055 7163 8105 8971
-Compilation failure 0.107 0.368 0.122 0.339 0.089 0.373
-OutOfMemory rate 0.304 0.241 0.471 0.369 0.448 0.312
-Table 4: Token usage and compilation failure rates. We report averages for input tokens and completion tokens per heuristic, and the percentage of compilation failures.
-Figure 3: Variance per domain for the tsr5 GPT-4.1 unrefined configuration over three independent runs, each limited to 600 seconds and 8 GB of memory per instance.
-Total coverages were 282, 282, and 276, indicating stable performance.
-All runs outperform the domain-independent SOTA.
-For comparison, we also include results from a run without the strategizing phase.
-Summary
-The ablation results suggest that refinement has minor positive effects, primarily due to increasing the diversity of heuristic candidates.
-The strategizing step, however, does not yield measurable benefit.
-These findings support the use of simple multi-sampling strategies to increase coverage, while questioning the need for more elaborate prompting constructs in heuristic generation.
+As a qualitative measure of the generated heuristics we report the head-to-head time and expanded nodes for solution comparison of sp-10 with GPT-4.1 vs.
+the best performing $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, which can be seen in Figure 5.Comparisons with additional models can be seen in Appendix.
+We conducted ablation studies to evaluate the contribution of three components in our heuristic generation pipeline:(i) the provision of the specific instance (ii) the model and fallback choice (iii) the reasoning effort requested from reasoning models.
+The study aimed to determine their impact on heuristic quality, generation time, and planning performance, as well as the variance in our best configuration.
+Impact of Instance-Specificity
+We measure the effect of providing the instance JSON to the model before generation, containing the initial state and the goal.
+This comes at the steep cost of requires generating additional heuristics per instance, as opposed to domain-specific-instance-general which can be generated once then ran on all future instances.
+Intuitively, IS may have the following benefits:
+  1. 1.
+    The heuristic often contains multiple penalizing elements which are combined with guessed coefficients.
+Knowing the details of the instance gives the model a better idea on how to tune these coefficients.
+  2. 2.
+    It tells the model how significant each part of the problem is.
+Consider a simple instance of Pacman where the ghosts are trapped.
+The model can avoid spending generation- and run-time on a complex subroutine to quantify the risk from the ghosts.
+  3. 3.
+    It tells the model the scale of the problem which can provide guidance on efficiency constraints.
+We hypothesize that there exists a tradeoff between the precision of the heuristic estimates and the computation time, and further that for larger problems it can be worthwhile for the heuristic to sacrifice precision for efficiency.
+For example, on TPP, Anthropic’s models implement path-finding algorithms like Dijkstra, Prim and Floyd-Warshall, about half the time.
+Although this is a good basis for a “smart” heuristic, it causes a timeout within a few thousand expansions for even medium problems.
+Meanwhile even Manhattan Distance heuristics achieves 16/20.
+  4. 4.
+    At times, a solution state is obvious from the initial state (e.g, FO-Counters).
+In these cases showing the model the instance details allows it to write a heuristic to guide efficiently toward that solution rather than abstractly guiding toward all solutions.
+Empirically, IS leads to a decrease in overall coverage (Table 2) with most domains varying slightly but a few increasing or decreasing significantly.
+IS did not cause heuristic generation time to increase [^9] Interestingly, despite adding tokens and theoretically adding complexity which can increase output tokens, adding the instance slightly decreased generation time.
+(P=2e-5<<0.05).
+We do not have an explanation for this fact..
+IS moderately increased compilation errors (see Table  3), and if this effect is normalized against, IS-heuristics have a slightly higher solving rate.
+We hypothesize that GPT-4.1 is insufficiently strong to make use of this info and that for a stronger model it may be worth revisiting.
+FC US SP-10
+Domain $D$ IS $D$ IS $D$ IS
+Block Grouping (20) 16 9 19 10 19 10
+Counters (20) 10 8 10 11 10 11
+Delivery (20) 16 10 20 18 20 18
+Drone (20) 15 15 16 16 16 16
+Expedition (20) 3 1 3 4 3 4
+Farming (20) 18 19 19 20 19 20
+FO-Counters (20) 4 8 4 14 4 14
+FO-Farming (20) 20 18 20 20 20 20
+FO-Sailing (20) 20 10 20 19 20 20
+Hydropower (20) 8 7 12 9 12 9
+Market Trader (20) 20 1 20 3 20 3
+Pathways (20) 0 1 1 1 1 1
+Plant Watering (20) 9 12 20 18 20 18
+Rover (20) 3 2 4 3 4 3
+Sailing (20) 11 14 20 20 20 20
+Settlers (20) 1 2 1 3 3 3
+TPP (20) 13 11 16 13 16 15
+Zenotravel (20) 20 14 20 20 20 20
+$\sum$ (360) 207 162 245 222 247 225
+Twin Prime (20) 18 20 18 20 18 20
+Pacman (20) 9 13 9 13 16 14
+$\sum$ (400) 234 195 272 255 281 259
+Table 2: Coverage results for GPT-4.1 with domain-dependence only ($D$), our standard configuration, against GPT-4.1 with Instance-Specificity (IS), reporting each on fc, us, and their best configuration, SP-10.
+For most domains GPT-4.1 fails to leverage the additional information, however we highlight FO-Counters and Twin Prime as domains where it lead to the most improvement.
+Impact of model and fallback choice
+We tested 6 models: GPT-4.1, GPT-4.1-mini, GPT-5.1, GPT-5-mini, Claude Sonnet 4.5, and Claude Haiku 4.5, each with each of our fallback options including fc, us, sp-5, and sp-10.
+For the full table see Appendix.
+Our observations are as follows:
+  1. 1.
+    For non-reasoning models, under about every configuration, a model’s larger variant is preferred.
+Since they are well-capable of generating and running many heuristics within the allotted time, SP10SP-10 is preferred.
+  2. 2.
+    Reasoning models are more competitive between sizes. They take longer to generate and $US$ performs best, allowing them the time they need.
+  3. 3.
+    Anecdotally, when reasoning models fail it can often be because they over-complicate the problem, taking longer to generate, and whether they end up with a better heuristic or an unusable one is a coin toss.
+That said, since we can re-sample this risk may be worth taking if the samplings are sufficiently independent.
+Impact of Reasoning effort
+We tested our method on the same model, GPT-5.1 with two configurations: reasoning_effort=’low’ and reasoning_effort=’high’ (for brevity denoted ’GPT-5.1’ and ’GPT-5.1=high’).
+We found that although increasing reasoning_effortreasoning\_effort slightly increased heuristic efficiency and quality, and better-than-halved compilation error rate, when we account for generation time the reduced number of heuristics we generate in time does not make this trade worthwhile.
+With better engineering one could generate many completions simultaneously, which may turn this trade worthwhile.
+We leave this direction for future work.
+For further graphs see appendix.
+GPT Claude
+Metric 4.1 4.1 (IS) 4.1 mini 5.1 5 mini 5.1 high 5 m-h Sonnet 4.5 Haiku 4.5
+Input tok. 3725 8102 3725 3724 3724 3724 3724 4761 4761
+Output tok. 1510 1695 1245 3253 2298 18075 14485 2122 1490
+Reasoning tok. [^10] For GPT, subset of output tokens. For Claude, they are not provided by the API - - - 1282 589 15510 11613 - -
+Cost (USD) 0.02 0.03 0.003 0.037 0.006 0.185 0.03 0.046 0.012
+Generation (sec.) 35.8 27.5 19.4 44.9 44.5 235.1 202.3 33.6 12.5
+Run (sec.) 24.7 22.1 19.2 24.2 18.0 21.4 36.7 43.4 22.2
+Compl. error % 0.300 0.430 0.215 0.150 0.405 0.070 0.205 0.195 0.165
+OOM rate % 0.304 0.241 0.388 0.382 0.280 0.340 0.288 0.391 0.451
+Runtime error % 0.027 0.020 0.035 0.032 0.028 0.035 0.032 0.016 0.032
+Timeout % 0.015 0.014 0.004 0.007 0.009 0.007 0.031 0.032 0.002
+Success % 0.355 0.295 0.358 0.430 0.278 0.548 0.444 0.365 0.351
+Table 3: Average token usage, cost, latency, and execution outcomes for different GPT and Claude models, independent of downstream applications.
+Models marked with are reasoning models.
+Excepting IS, input tokens across models depend only on the tokenizer.
+Costs do not account for token caching discounts.
+Figure 4: As a variance analysis we generated 40 heuristics with GPT-4.1 in each domain, and performed a Monte-Carlo simulation of our algorithm for 1000 iterations of sampling heuristic order, taking the coverage per-domain and overall each time.
+The boxes represent the first and third quartiles, the thick line the median, and the whiskers the 1st and 99th percentiles.
+Although limited, the analysis shows the method to be highly consistent for most domains.
+The ablation results suggest that non-reasoning models don’t significantly benefit from IS.
+The best setting of our method is to use a non-reasoning model to generate non-IS heuristics and run them as SelfPortfolio-10 (276 with GPT-4.1).
+If time is allowed ahead of receiving the instances, the best setting becomes instead to generate ahead of time as many heuristics as possible and run SP-N or US on them (The two fallbacks are approximately equivalent when used on pregenrated heuristics due to near-zero timeout rate).
 Discussion
 As shown in Table 1, LLM-based heuristics achieve state-of-the-art performance across the evaluated domains.
-They are particularly effective in challenging benchmarks such as Sailing and Zenotravel, which are known to be difficult for domain-independent planners.
-The best-performing LLM-based heuristics were produced by GPT-4.1, with Sonnet 3.5 yielding comparable results.
+They are particularly effective in challenging benchmarks such as (FO-) Sailing and Zenotravel, which are known to be difficult for domain-independent planners.
+The best-performing LLM-based heuristics were produced by GPT-5.1 and it’s mini variant with high reasoning effort, however when accounting for generation time and cost, GPT-4.1 emerges as a clear winner.
 Nonetheless, certain domains remain challenging for LLMs, likely due to their sensitivity to non-descriptive variable names and opaque goal specifications, as seen in Pathways and Settlers.
@@
 This suggests that departing from pddl-based input to a direct successor-generator representation may, in some cases, degrade performance, allowing us to reject the hypothesis that our representation alone accounts for the observed improvements.
-Costs and Efficiency While concerns about the resource efficiency of LLMs have been raised (Stojkovic et al.
-2024), energy efficiency is beyond the scope of this work.
-Nonetheless, our method is amenable to efficient deployment: the first two phases are domain-specific rather than instance-specific and can be executed once per domain, substantially reducing energy and cost.
-On average, solving an input using TSR5 with GPT-4.1 requires 56k input tokens and 8.8k output tokens (Table 4), corresponding to a pre-caching cost of approximately 0.18 USD.
+Costs and Efficiency While concerns about the resource efficiency of LLMs have been raised (stojkovic2024towards), energy efficiency is beyond the scope of this work.
+Nonetheless, our method is amenable to efficient deployment: generating heuristics can be done ahead of time one per domain then reused on every instance, substantially reducing energy and cost even if a powerful model is used.
+Additionally, the resulting heuristics tend to be very efficient to run, and succeed or fail much quicker than domain-independent heuristics (see Appendix.)
 Conclusion